OpenStudy (anonymous):

Determine the range for: g(x)=1/2(6+x)(4+x) I do not know how to do domain and range

OpenStudy (zzr0ck3r):

can you find the inverse?

OpenStudy (anonymous):

i am not entirely sure how to do that either. :/ math is really not my best.

OpenStudy (zzr0ck3r):

\[\frac{1}{2(6+x)(4+x)}=g(x)\]

OpenStudy (zzr0ck3r):

is that what that says?

OpenStudy (anonymous):

hold on, i'll attach a picture

OpenStudy (zzr0ck3r):

this is porabolla opening up, So its range will be above the vertex vertex = -b/2a x=-5 plug that in y=-1/2 so its all y>=-1/2

OpenStudy (anonymous):

ok, thank you zzr0ck3r